OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? b + 14 = 25 someone please help me ok

OpenStudy (anonymous):

not add taken away. just put -14 on both sides

OpenStudy (anonymous):

so b+14-14=25-14 b=11